### v1

# Lower tensor-to-scalar ratio as possible signature of modified gravity

Abstract
This paper simplifies the induced four-dimensional gravitational equations originating from a five-dimensional bulk within the framework of Nash’s embeddings, incorporating them into a well-known $\mu-\Sigma$ modified gravity (MG) parametrization. By leveraging data from Planck Public Release 4 (PR4), BICEP/Keck Array 2018, Planck cosmic microwave background lensing, and baryon acoustic oscillation observations, we establish a stringent lower limit for the tensor-to-scalar ratio parameter: $r<0.0303$ at a confidence level (CL) of 95%. This finding suggests the presence of extrinsic dynamics influencing standard four-dimensional cosmology. Notably, this limit surpasses those typically obtained through Bayesian analysis using Markov Chain Monte Carlo (MCMC) techniques, which yield $r<0.038$, or through the frequentist profile likelihood method, which yields $r<0.037$ at 95% CL.
preprint: APS/123-QED
I Introduction
As largely accepted, inflation became one of the cornerstone of modern cosmology. It not only solves the flatness and horizon problems, but also states the quantum seeds of cosmological fluctuations that eventually drove the universe to evolve [1, 2, 3, 4, 5, 6, 7, 8, 9] throughout a brief and rapid exponential expansion of the universe right after Big Bang. This inflationary period is thought to have smoothed out the early universe’s irregularities and laid the foundation for the large-scale structure we observe today. It is reinforced by the anisotropy seen in Cosmic Microwave Background (CMB) observations [10]. While the precise mechanisms fueling inflation remain elusive, numerous competing theoretical models have emerged [11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21] , each vying to explain this fundamental cosmic process. However, despite the ongoing debate, the overarching framework consistently yields predictions that align remarkably well with cosmological observations (see [22, 23] for a review).
A significant portion of scientific inquiry into the origins of the universe revolves around scrutinizing and characterizing the statistical properties of primordial density perturbations, particularly through the analysis of the statistical two-point function. Empirical evidence suggests that these fluctuations adhere closely to Gaussian distribution and exhibit near-scale invariance. Consequently, we can effectively capture their statistical behavior using a power-law power spectrum governed by two crucial parameters. These parameters, integral to the $\Lambda$CDM standard cosmological model, are the scalar amplitude $A_{s}$, representing the perturbation amplitude, and the spectral tilt $n_{s}$, governing the scale dependency of the density perturbation power spectrum. Remarkably, even with just these two parameters, we can glean insights into certain facets of inflationary dynamics today, shedding light on the energy scales pivotal during the early epochs of the universe’s evolution. Statistical analyses from canonical inflationary models constrain $n_{s}=0.9649\pm 0.0044$ and $\ln(10^{10}A_{s})=3.044\pm 0.014$ at the 68% confidence level (CL) using Planck-CMB data [24]. Conversely, analysis from ACT-CMB indicates agreement with a Harrison-Zel’dovich primordial spectrum, where $n_{s}=1.009\pm 0.015$ [25]. This introduces tension with Planck-CMB measurements. For further discussions on this topic, see also [26, 27, 28].
[…]
In this paper, we propose a possibility of the extrinsic curvature works as an inflaton field. Differently from traditional extra-dimensional models [32, 33, 34, 35], the extrinsic curvature is thought as the orthogonal component of the gravitational field besides the metric $g_{\mu\nu}$ as its tangent component. Our leitmotiv focus on how a dynamic embedding should affect the physical universe and may provide new insights on the current problems in cosmology. As a main character in this framework, the extrinsic curvature should not be restricted to the analysis of the background geometry but it should also include its perturbative dynamics [36, 37, 38]. To do so, we adopt Nash-Greene theorem [39, 40] that provides a general structure for embedding between non-Riemannian geometries [41, 42, 43, 44, 45, 46, 47, 48, 49, 50]. It worth noting that, we do not use specific conditions on the extrinsic curvature such as Israel-Darmois-Lanczos (IDL) condition [51]. Commonly used in Randall-Sundrum models [33, 34], IDL condition replaces the extrinsic curvature by an algebraic relation with the energy-momentum tensor. It was showed it belongs to a very special case [52, 42], and, in general, IDL condition can be completely removed or replaced [53].
The paper is organized in sections. The second section verses on the essentials of embeddings as a mathematical background and summarises our cosmological model with both background and perturbed cosmological equations as well as the effective fluid approach necessary to develop a more realistic model to compare with observations. In the third section, we present a simpler version of the model in the form of a Modified Gravity (MG) framework. The fourth section is devoted to the analysis of extrinsic curvature as an inflaton field in which we investigate the slow-roll conditions for the extrinsic inflation potential. In sections V and IV, we examine the model in contrast with cosmological data with the latest NPIPE Planck DR4 likelihoods [54, 55], the BICEP2/Keck collaboration [56] and a junction of Large-scale structure (LSS) catalogue with 6dF Galaxy Survey [57], the Seventh Data Release of SDSS Main Galaxy Sample (SDSS DR7 MGS) [58] and clustering measurements of the Extended Baryon Oscillation Spectroscopic Survey (eBOSS) associated with the SDSS’s Sixteenth Data Release [59]. The related joint likelihood analysis is computed with MGCAMB-II [60] patch by means of Cobaya [61] sampler. In the final section, we conclude with our remarks and prospects.
II Embeddings as a theoretical background
We present the main results from previous works [41, 42, 43, 44, 49]. Once a general arbitrary D-dimensional case is possible [41, 42], for comparison purposes in the recent literature, we define our model in five dimensions that has a five-dimensional bulk $V_{5}$ with a embedded four-dimensional geometry $V_{4}$.
A gravitational action $S$ is defined in such a way
$S=-\frac{1}{2\kappa^{2}_{5}}\int\sqrt{|\mathcal{G}|}{}^{5}\mathcal{R}d^{5}x-\int\sqrt{|\mathcal{G}|}\mathcal{L}^{*}_{m}d^{5}x\;,$ (1)
where $\kappa^{2}_{5}$ is a fundamental energy scale on the embedded space, the curly ${}^{5}\mathcal{R}$ curvature means the five-dimensional Ricci scalar, and the matter Lagrangian $\mathcal{L}^{*}_{m}$ denotes the confined matter fields on a four-dimensional embedded space-time.
The non-perturbed extrinsic curvature $\bar{k}_{\mu\nu}$ is defined as [62],
$k_{\mu\nu}=-\mathcal{X}^{A}_{,\mu}\;\eta^{B}_{,\nu}\mathcal{G}_{AB}\;.$ (2)
Eq.(2) shows the projection of the variation of normal unitary vector $\eta^{B}$ onto the tangent plane. The embedding coordinate $\mathcal{X}$ defines a regular local map $\mathcal{X}:V_{4}\rightarrow V_{5}$. The bulk metric $\mathcal{G}_{AB}$ is defined as
$\mathcal{G}_{AB}$ $=$ $\left(\begin{array}[]{cc}g_{\mu\nu}&0\\ 0&1\\ \end{array}\right)\;.$ (5)
Concerning notation, capital Latin indices run from 1 to 5. Overbar terms are background quantities unless. Small case Latin indices refer to the only one extra dimension considered. All Greek indices refer to the embedded space-time counting from 1 to 4.
In this paper, we use Nash theorem of 1956 [39]. The seminal result of this theorem shows how to produce orthogonal perturbations from the background metric $\bar{g}_{\mu\nu}$ given by
$\bar{k}_{\mu\nu}=-\frac{1}{2}\frac{\partial\bar{g}_{\mu\nu}}{\partial y}\;.$ (6)
where $y$ is an arbitrary spatial coordinate. As a consequence of Eq.(6), we have how new geometries $g_{\mu\nu}$ are generated from small perturbations on the background metric increments $\delta g_{\mu\nu}$ as
$g_{\mu\nu}=\;\bar{g}_{\mu\nu}+\delta g_{\mu\nu}\;.$ (7)
Moreover, one obtains the perturbed extrinsic curvature $k_{\mu\nu}$ as
$k_{\mu\nu}=\;\bar{k}_{\mu\nu}+\delta k_{\mu\nu}\;,$ (8)
where $\delta k_{\mu\nu}=-2\delta y\;g^{\sigma\rho}\;k_{\mu\sigma}k_{\nu\rho}$. In principle, this is a continuous process to an any arbitrary perturbation increments of superior orders of $\delta g_{\mu\nu}$ and $\delta k_{\mu\nu}$. For our purposes, we keep the perturbations at linear order as show in Eq.(7) and (8).
The increments $\delta g_{\mu\nu}$ and $\delta k_{\mu\nu}$ are the Nash’s fluctuations later applied to non-Riemannian metrics by Greene [40]. The deformation formula Eq.(6) is also a pivot element to obtain solutions of the Gauss and Codazzi equations
${}^{5}{\cal R}_{ABCD}\mathcal{Z}^{A}{}_{,\alpha}\mathcal{Z}^{B}{}_{,\beta}\mathcal{Z}^{C}{}_{,\gamma}\mathcal{Z}^{D}{}_{,\delta}=\bar{R}_{\alpha\beta\gamma\delta}\;+$
$\hskip 113.81102pt+\;(\bar{k}_{\alpha\gamma}\;\bar{k}_{\beta\delta}\!-\!\;\bar{k}_{\alpha\delta}\;\bar{k}_{\beta\gamma})\!\;,$ (9)
${}^{5}{\cal R}_{ABCD}\mathcal{Z}^{A}{}_{,\alpha}\mathcal{Z}^{B}{}_{,\beta}\mathcal{Z}^{C}{}_{,\gamma}\eta^{D}=\;\bar{k}_{\alpha[\beta;\gamma]}\;,$ (10)
where ${}^{5}{\cal R}_{ABCD}$ is the five-dimensional Riemann tensor and $\bar{R}_{\alpha\beta\gamma\delta}$ is the background four-dimensional Riemann tensor. The perturbed embedding coordinate $\mathcal{Z}^{A}_{,\mu}$ is defined as $\mathcal{Z}^{A}_{,\mu}=\mathcal{X}^{A}_{,\mu}+\delta y\;\eta^{A}_{,\mu}\;$. The normal vector $\eta^{A}$ is invariant under perturbations, i.e., $\eta^{A}=\bar{\eta}^{A}$. The semicolon in Eq.(10) denotes covariant derivative with respect to the metric and the brackets apply the covariant derivatives to the adjoining indices.
II.1 Background cosmology
The background cosmology is defined as usual by means of the line element of Friedmann-Lemaître-Robertson-Walker (FLRW) four-dimensional metric as
$ds^{2}=-dt^{2}+a^{2}\left(dr^{2}+r^{2}d\theta^{2}+r^{2}\sin^{2}\theta d\phi^{2}\right)\;,$ (11)
where the scale factor is denoted by $a\equiv a(t)$ and $t$ is the physical time. Once the embedding is properly set, we impose that the bulk geometry is a solution of Einstein’s equations. Thus, from Eq.(9) and  Eq.(10), one obtains the non-perturbed field equations as
$\bar{G}_{\mu\nu}-\bar{Q}_{\mu\nu}$ $=$ $-8\pi G\bar{T}_{\mu\nu}\;,$ (12)
$\bar{k}_{\mu[\nu;\rho]}$ $=$ $0\;,$ (13)
where $\bar{T}_{\mu\nu}$ is the non-perturbed energy-momentum tensor of the confined perfect fluid and $G$ is the gravitational Newtonian constant. The background tensors $\bar{G}_{\mu\nu}$ and $\bar{Q}_{\mu\nu}$ represent the four dimensional Einstein tensor and the extrinsic deformation tensor, respectively. For an arbitrary D-dimensional case, see the detail derivation of Eq.(12) and Eq.(13[42] .
We can define the non-perturbed $\bar{Q}_{\mu\nu}$ in Eq.(12) is given by
$\bar{Q}_{\mu\nu}=\bar{k}^{\rho}_{\mu}\bar{k}_{\rho\nu}-\bar{k}_{\mu\nu}h-\frac{1}{2}\left(K^{2}-h^{2}\right)\bar{g}_{\mu\nu}\;,$ (14)
where $h^{2}=h\!\cdot\!h$ denotes the mean curvature, being $h=\;\bar{g}^{\mu\nu}\;\bar{k}_{\mu\nu}$. The Gaussian curvature is denoted by $K^{2}=\bar{k}^{\mu\nu}\bar{k}_{\mu\nu}$. A direct consequence of the previous definition in Eq.(14) is that the deformation tensor $\bar{Q}_{\mu\nu}$ is a conserved quantity such as
$\bar{Q}_{\mu\nu;\mu}=0\;.$ (15)
As an FLRW solution, the trace of Codazzi equation of Eq.(13) and Eq.(14) were found as [42]
$\bar{k}_{ij}=\frac{b}{a^{2}}\bar{g}_{ij},\;\;i,j=1,2,3,$ (16)
$\bar{k}_{44}=\frac{-1}{\dot{a}}\frac{d}{dt}\frac{b}{a}$ (17)
$\bar{k}_{44}=-\frac{b}{a^{2}}\left(\frac{B}{H}-1\right)\;,$ (18)
$K^{2}=\frac{b^{2}}{a^{4}}\left(\frac{B^{2}}{H^{2}}-2\frac{B}{H}+4\right),\;\,h=\frac{b}{a^{2}}\left(\frac{B}{H}+2\right),$ (19)
$\bar{Q}_{ij}=\frac{b^{2}}{a^{4}}\left(2\frac{B}{H}-1\right)\bar{g}_{ij},\;\bar{Q}_{44}=-\frac{3b^{2}}{a^{4}},$ (20)
$\bar{Q}=-(K^{2}-h^{2})=\frac{6b^{2}}{a^{4}}\frac{B}{H}\;,$ (21)
where $\bar{Q}$ denotes the deformation scalar defined in a standard way, i.e., by the contraction $\bar{g}^{\mu\nu}\bar{Q}_{\mu\nu}=\bar{Q}$. One important definition is the evolution of the bending function $b(t)=b\equiv b(t)=k_{11}=k_{22}=k_{33}$ driven by extrinsic geometry. Thus, we defined $B=B(t)\equiv\frac{\dot{b}}{b}=(db/dt)/b$ as a copy of the Hubble parameter $H\equiv H(t)=\frac{\dot{a}}{a}=(da/dt)/a$. The $B(t)$ function has the same unit as $H$.
As shown in [42, 44], one obtains
$b(t)=b_{0}a(t)^{\beta_{0}}\;,$ (22)
where the term $b_{0}$ is the current value of the bending function and $\beta_{0}$ is an integration constant. As we are going to show, it will be associated with the fluid parameter $w$. Accordingly, the Friedmann equation in terms of redshift can be written in a form
$\left(\frac{H}{H_{0}}\right)^{2}=\Omega_{m(0)}(1+z)^{3}+\Omega_{rad0}(1+z)^{4}$
$\hskip 113.81102pt+\Omega_{ext(0)}(1+z)^{4-2\beta_{0}},$ (23)
where $\Omega_{m(0)}$ denotes the current cosmological parameter for matter density content. The radiation content is denoted by $\Omega_{rad(0)}=\Omega_{m0}z_{eq}$, wherein the equivalence number for the expansion factor $a_{eq}$ is
$a_{eq}=\frac{1}{1+z_{eq}}=\frac{1}{(1+2.5\times 10^{4}\Omega_{m(0)}h^{2}(T_{cmb}/2.7)^{-4})}\;,$ (24)
where $z_{eq}$ is the equivalence redshift. The CMB temperature is adopted for the value $T_{cmb}=2.7255K$ and the Hubble factor $h=0.67$ [10]. The term $\Omega_{ext(0)}$ stands for the density parameter associated with the extrinsic curvature. For a flat universe, $\Omega_{ext(0)}=1-\Omega_{m(0)}-\Omega_{rad(0)}$. Moreover, the current extrinsic cosmological parameter $\Omega_{ext(0)}$ is defined as
$\Omega_{ext(0)}=\frac{8\pi G}{3H_{0}^{2}}\rho_{ext(0)}\equiv\frac{b_{0}^{2}}{H_{0}^{2}a_{0}^{\beta_{0}}}\;,$ (25)
where $a_{0}$ sets the current value of the scale factor and $\rho_{ext(0)}\equiv\frac{3}{8\pi G}b^{2}_{0}$ denotes the current extrinsic density.
II.2 Perturbations in Conformal Newtonian gauge
The perturbed equations are necessary for the right estimation of cosmological parameters. In this framework, the relevant modifications of field equations under cosmological perturbations are applied to LHS of Eq.(12) with the presence of $\bar{Q}_{\mu\nu}$, and $\bar{G}_{\mu\nu}$ and $\bar{T}_{\mu\nu}$ tensors are treated in a very standard fashion as shown in Ref. [44]. Thus, a fluid with pressure $P$ and density $\bar{\rho}$, the perturbed components of the stress-tensor energy $\delta T_{\mu\nu}$ are given as
$\delta T^{i}_{j}$ $=$ $\delta p\;\delta^{i}_{j}+\Sigma_{j}^{i}\;,$ (26)
$\delta T^{4}_{4}$ $=$ $-\delta\rho\;,$ (27)
$\delta T^{4}_{i}$ $=$ $\frac{1}{a}(\bar{\rho}+P)\delta u_{\parallel i}\;,$ (28)
where $\delta u_{\parallel i}$ is the tangent velocity potential of the fluid. The anisotropic stress tensor is defined as $\Sigma_{j}^{i}=T^{i}_{j}-\delta^{i}_{j}T^{k}_{k}/3$. Moreover, the only relevant field equations that propagate cosmological perturbations are given by
$\delta G^{\mu}_{\nu}-\delta Q^{\mu}_{\nu}=-8\pi G\delta T^{\mu}_{\nu}\;.$ (29)
In addition, we have to determine the perturbed extrinsic terms given by $\delta Q_{\mu\nu}$. Using the Nash relation of Eq.(6) and Eq.(8), one obtains
$\delta{k}_{\mu\nu}=\;\bar{g}^{\sigma\rho}\;\bar{k}_{\mu\sigma}\;\delta g_{\nu\rho}\;.$ (30)
This is a pivot result since it shows how the effects of the extrinsic quantities are projected onto the perturbed four embedded space-time and on how Nash flow of Eq.(6) is connected to cosmological perturbations. Hence, the perturbations of the deformation tensor $\bar{Q}_{\mu\nu}$ are given in a form
$\delta Q_{\mu\nu}=-\frac{3}{2}(K^{2}-h^{2})\delta g_{\mu\nu}\;.$ (31)
Alternatively, we construct a relation of the equations in a fluid approach writing the gravi-tensor equation in a general form as
$G_{\mu\nu}=-8\pi GT^{total}_{\mu\nu}\;,$ (32)
where the tensors are written by the composition of their background and perturbed components. The related Einstein tensor is written as $G_{\mu\nu}=\bar{G}_{\mu\nu}+\delta G_{\mu\nu}$, $T^{total}_{\mu\nu}=T_{\mu\nu}+T^{ex}_{\mu\nu}$, where $T_{\mu\nu}=\bar{T}_{\mu\nu}+\delta T_{\mu\nu}$ and $T^{ex}_{\mu\nu}=\bar{T}^{ex}_{\mu\nu}+\delta T^{ex}_{\mu\nu}$.
In the conformal Newtonian gauge, the FLRW metric is given by
$ds^{2}=a^{2}[-(1+2\Psi)d\tau^{2}+(1-2\Phi)dx^{i}dx_{i}]\;,$ (33)
where $\Psi=\Psi(\vec{x},\tau)$ and $\Phi=\Phi(\vec{x},\tau)$ denote the Newtonian potential and the Newtonian curvature in conformal time $\tau$ that is defined as $d\tau=dt/a(t)$. Hence, we determine the confined matter fields represented by the non-perturbed stress energy tensor $\bar{T}_{\mu\nu}$ in a co-moving fluid such as
$\bar{T}_{\mu\nu}=\left(\bar{\rho}+P\right)U_{\mu}U_{\nu}+P\bar{g}_{\mu\nu}\;;\;U_{\mu}=\delta^{4}_{\mu}\;,$ (34)
where $U_{\mu}$ denotes the co-moving velocity. The related conservation equation is given by
$\bar{\rho}+3H\left(\bar{\rho}+P\right)=0\;.$ (35)
From the perturbed conservation equation $\delta{T}_{\mu\nu;\nu}=0$, the evolution equations can be obtained for the “contrast” matter density $\delta_{m}$ and fluid velocity $\theta$ such as
$\dot{\delta_{m}}=-(1+w)(\theta-3\dot{\Phi})-3H(c_{s}^{2}-w)\delta_{m}\;,$ (37)
$\dot{\theta}=-H(1-3w)\theta-\frac{\dot{w}}{1+w}\theta+\frac{c_{s}^{2}}{1+w}k^{2}\delta_{m}+$
$\hskip 142.26378pt-k^{2}\sigma+k^{2}\Psi\;,$
where $\theta=ik^{j}u_{j}$, $w$ is the fluid parameter $w=\frac{P}{\bar{\rho}}$, $c^{2}_{s}$ is the sound velocity defined as $c^{2}_{s}=\frac{\delta P}{\delta\rho}$ and $\sigma$ is the anisotropic stress. The dot symbol denotes the ordinary derivative with respect to conformal time $\tau$.
To avoid divergences when Equation of State (EoS) crosses $w=-1$, eventually, the scalar velocity $V=(1+w)\theta$ should be defined [63, 64]. Hence, we have the equations
$\delta^{\prime}_{m}=3(1+w)\Phi^{\prime}-\frac{V}{a^{2}H}-\frac{3}{a}\left(\frac{\delta P}{\bar{\rho}}-w\delta_{m}\right)\;,$ (39)
$V^{\prime}=-(1-3w)\frac{V}{a}+\frac{k^{2}}{a^{2}H}\frac{\delta P}{\bar{\rho}}+(1+w)\frac{k^{2}}{a^{2}H}\Psi+$
$\hskip 128.0374pt-\frac{k^{2}}{a^{2}H}(1+w)\sigma,$
where the prime symbol ${}^{\prime}$ denotes the ordinary derivative with respect to scale factor as ${}^{\prime}=\frac{d}{da}$.
For the induced extrinsic part, we have that the term $8\pi G\bar{T}^{ext}_{\mu\nu}\doteq\bar{Q}_{\mu\nu}$ which is written as copy of a perfect fluid as
$-8\pi G\bar{T}^{ext}_{\mu\nu}=(\bar{p}_{ext}+\bar{\rho}_{ext})U_{\mu}U_{\nu}+\bar{p}_{ext}\;\bar{g}_{\mu\nu},\;\;\;U_{\mu}=\delta_{\mu}^{4}\;,$ (40)
where $U_{\mu}$ is the co-moving four-velocity. Since $T_{\mu\nu;\nu}=0$ and $T^{ext}_{\mu\nu;\nu}=0$ (as a consequence of Eq.(15)), $T^{total}_{\mu\nu}$ is conserved, accordingly. Hence, the conservation equation for extrinsic quantities is given by
$\frac{d\bar{\rho}_{ext}}{dt}+3H\left(\bar{\rho}_{ext}+\bar{p}_{ext}\right)=0\;,$ (41)
where $\bar{\rho}_{ext}$ and $\bar{p}_{ext}$ denote the non-perturbed extrinsic density and extrinsic pressure, respectively. The time-time and space-time components of $T^{ext}_{\mu\nu}$ can be set as
$-8\pi GT^{ext}_{44}=\bar{T}^{ext}_{44}+\delta T^{ext}_{44}=\bar{Q}_{44}+\delta Q_{44}\;,$ (42)
$-8\pi GT^{ext}_{i4}=\bar{T}^{ext}_{i4}+\delta T^{ext}_{i4}=\bar{Q}_{i4}+\delta Q_{i4}=0\;.$ (43)
The modified Friedman equation from is written as
$H^{2}=\frac{8}{3}\pi G\left(\bar{\rho}_{m}+\bar{\rho}_{rad}+\bar{\rho}_{ext}\right)\;\;,$ (44)
where $\bar{\rho}_{ext}(a)$ is given by
$\bar{\rho}_{ext}(a)=\bar{\rho}_{ext}(0)a^{2\beta_{0}-4}\;,$ (45)
with $\bar{\rho}_{ext}(0)=\frac{3}{8\pi G}b_{0}^{2}$. Once $\bar{\rho}_{ext}(a)$ is already determined, the “extrinsic” pressure can be calculated using Eqs.(41) and (45) to obtain
$\bar{p}_{ext}(a)=\frac{1}{3}\left(1-2\beta_{0}\right)\bar{\rho}_{ex}(0)a^{2\beta_{0}-4}\;.$ (46)
Concerning the Null Energy Conditions (NEC), for positive values of $\beta_{0}$, NEC is satisfied. If $\beta_{0}=2$, the extra term in the modified Friedmann equation in Eq.(II.1) mimics a cosmological constant. If $\beta_{0}>2$, the effective behavior becomes phantom-like pattern. However, the condition $\beta_{0}>1$ is a must if the universe must accelerate (apart from being consistent with NEC). Thus, we define an effective EoS with an “extrinsic fluid” parameter $w_{ext}$ by the definition $w_{ext}=\frac{\bar{p}_{ext}}{\bar{\rho}_{ext}}$ to obtain
$w_{ext}=-1+\frac{1}{3}\left(4-2\beta_{0}\right)\;.$ (47)
which when
$\beta_{0}=2-\frac{3}{2}\left(1+w\right),$ (48)
one has the fluid correspondence $w_{ext}=w$. This allows us to express all the relevant quantities in a fluid approach. For instance, the dimensionless Hubble parameter $E(z)=\frac{H(z)}{H_{0}}$ is written as
$E^{2}(z)=\Omega_{m(0)}(1+z)^{3}+\Omega_{rad(0)}(1+z)^{4}+$
$\hskip 113.81102pt+\Omega_{ext(0)}(1+z)^{3(1+w)}.$ (49)
Just for referencing purposes, we name this model as $\beta$-model.
From Eq.(31), we calculate the set of non-zero components of $\delta Q_{\mu\nu}$ using the background relations in Eqs.(16), (19), (20) and (21). Then, one obtains
$\delta Q^{4}_{4}$ $=$ $\gamma_{0}a^{-(1+3w)}\Psi\delta^{4}_{4}\;,$ (50)
$\delta Q^{i}_{j}$ $=$ $\gamma_{0}a^{-(1+3w)}\Phi\delta^{i}_{j}\;,$ (51)
$\delta Q^{i}_{i}$ $=$ $3\gamma_{0}a^{-3(1+w)}\Phi\;,$ (52)
$\delta Q$ $=$ $\gamma_{0}(3\Psi-\Phi)a^{-3(1+w)}\;.$ (53)
Just for notation sake and using Eq.(25), in the foregoing set of equations $\gamma_{0}$ denotes
$\gamma_{0}=18b_{0}^{2}\beta_{0}=9H_{0}^{2}\Omega_{ext(0)}\gamma_{s}\;,$ (54)
where $\beta_{0}$ was simplified by the introduction of the $\gamma_{s}$ parameter that is regarded as a relic from extrinsic geometry to maintain the characteristic of the parameter $\gamma_{0}$. Moreover, it turns $\gamma_{0}$ in Eq.(54) independent of the nature of the fluid characterized by $w$ parameter, as expected. This is also necessary to maintain the RG correspondence that is achieved when extrinsic curvature vanishes, i.e., $\gamma_{s}\rightarrow 0$.
From Eq.(32) one writes the gauge invariant perturbed field equations in the Fourier k-space wave modes as
$k^{2}\Phi_{k}+3\mathcal{H}\left(\Phi^{{}^{\prime}}_{k}+\Psi_{k}\mathcal{H}\right)=-4\pi Ga^{2}\delta\rho_{k}+\chi(a)\Psi_{k},$ (55)
$\Phi^{{}^{\prime}}_{k}+\mathcal{H}\Psi_{k}=-4\pi Ga^{2}(\bar{\rho}+P)\frac{\theta}{k^{2}}\;,$ (56)
$\mathcal{D}_{k}+\frac{k^{2}}{3}(\Phi_{k}-\Psi_{k})=-\frac{4}{3}\pi Ga^{2}\delta\bar{P}-\frac{1}{2}a^{2}\delta Q^{i}_{i}\;,$ (57)
$k^{2}(\Phi_{k}-\Psi_{k})=12\pi Ga^{2}(\bar{\rho}+P)\sigma,$ (58)
where $\theta=ik^{j}\delta u_{\parallel j}$ denotes the divergence of fluid velocity in k-space, and $\mathcal{D}_{k}$ denotes $\mathcal{D}_{k}=\Phi^{{}^{\prime\prime}}_{k}+\mathcal{H}(2\Phi_{k}+\Psi_{k})^{\prime}+(\mathcal{H}^{2}+2\mathcal{H}^{\prime})\Psi_{k}$. The function $\chi(a)$ is expressed in terms of the cosmological parameters and reads
$\chi(a)=\frac{9}{2}\gamma_{s}\frac{H_{0}^{2}}{\Omega_{rad}(a)}\Omega_{rad}(0)\Omega_{ext}(a)\;.$ (59)
III Embedding as a modified gravity model
The use of the set of perturbed equations in Eqs.(55), (56, (57) and (58) is simplified after some algebra into the following set of equations
$k^{2}\Psi_{k}=-4\pi Ga^{2}\mu(a,k)\rho\Delta\;,$ (60)
$k^{2}(\Phi_{k}+\Psi_{k})=-8\pi Ga^{2}\Sigma(a,k)\rho\Delta\;,$ (61)
where $\rho\Delta=\bar{\rho}\delta+3\frac{\mathcal{H}}{k}(\bar{\rho}+P)\theta$. The set of Eqs.(60) and (61) are valid for all times. Thus, $\mu(a,k)$ and $\Sigma(a,k)$ functions can be written as
$\mu(a,k)=\frac{1}{1-\frac{\chi(a)}{k^{2}}}\left[1+\frac{3(\bar{\rho}+P)\sigma}{\rho\Delta}\right]\;,$ (62)
$\Sigma(a,k)=\frac{1}{2}\left[1+\mu(a,k)\left(1+\frac{\chi(a)}{k^{2}}\right)\right]\;.$ (63)
It is worth noting that the anisotropic stress is inserted in $\mu$ function. This is due to the fact that we do not define an extrinsic source, rather, the influence of extrinsic geometry is done over the matter content perturbing background geometry. Thus, the anisotropy, if considered, is a full matter-source.
Using the definition of the slip function $\gamma(a,k)=\frac{\Phi}{\Psi}$, from Eqs. (62) and (63), one easily obtains
$\Sigma(a,k)=\frac{1}{2}\mu(a,k)(1+\gamma(a,k))\;.$ (64)
When anisotropic stresses are neglected, one does not obtain the standard GR limit as $\Sigma(a,k)=\mu(a,k)$ and $\gamma(a,k)=1$. In this framework, the equality of Newtonian potentials $\Phi=\Psi$ is obtained in twofold conditions: anisotropic stress must vanish and the extrinsic term $\gamma_{s}\rightarrow 0$ in order to recover GR correspondence.
The set of basic equations are complete with the matter perturbation equations in Eqs.(39) and (39), and the evolution equation of $\Phi$. To obey solar constraints $\gamma_{s}$ in $\mu(a,k)$ function must comply with the condition
$\gamma_{s}<\frac{0.222k^{2}_{p}}{H_{0}^{2}\Omega_{ext(0)}}\;,$ (65)
[…]
IV Extrinsic curvature as an effective inflaton field
In previous publications [44, 48, 49], we explored some of the consequences of the Nash’s embedding in the context of dark energy problem. The appearance of the extrinsic energy density $\bar{\rho}_{ext}$ drives the universe to the late accelerated expansion and Friedmann equation is written in a shorter form as
$H^{2}=\frac{\kappa^{2}}{3}\left(\bar{\rho}+\bar{\rho}_{ext}\right)\;,$ (66)
with $\kappa=8\pi G$ and $\bar{\rho}=\bar{\rho}_{m}+\bar{\rho}_{rad}$. In terms of inflationary cosmology, such energy density $\bar{\rho}_{ext}$ should provide a response in a form of a scalar field potential energy density $V(\phi)$ generated by a spatially homogeneous “extrinsic” scalar field $\phi$ [50]. Thus, one defines a Lagrangian $\mathcal{L}_{\phi}$ such as
$\mathcal{L}_{\phi}=\frac{1}{2}\dot{\phi}^{2}-V(\phi)\;,$ (67)
where the time derivative is represented by the dot symbol. Immediately, one writes the related energy momentum tensor as
$T^{\phi}_{\mu\nu}=\partial_{\mu}\phi~{}\partial_{\nu}\phi+g_{\mu\nu}\left(\frac{1}{2}\dot{\phi}^{2}-V(\phi)\right)\;.$ (68)
From conservation of Eq.(68), the inflation dynamics is coupled to the background evolution, and we obtain the relations
$\bar{\rho}_{ext}=\frac{\dot{\phi}^{2}}{2}+V(\phi)\;,$ (69)
$\bar{p}_{ext}=\frac{\dot{\phi}^{2}}{2}-V(\phi)\;,$ (70)
where $\bar{\rho}_{ext}$ and $\bar{p}_{ext}$ denote the energy density and pressure generated by extrinsic curvature as a function of the extrinsic scalar field $\phi$.
As shown in ref [50], using Eqs.(45), (46), (47), (69) and (70), by direct integration, one obtains the potential $\phi(a)$ as
$\phi(a)=\sqrt{|3(1+w)|}~{}M_{pl}\ln{a}\;.$ (71)
Heron we denote the reduced Planck mass as $M_{pl}=\frac{1}{\sqrt{8\pi G}}=c=1$. The potential $V(\phi)$ is also obtained straightforwardly
$V(\phi)=V_{0}e^{-\alpha_{0}\phi}\;,$ (72)
where $\alpha_{0}=\sqrt{|3(1+w)|}$. This kind of exponential model is well-known in literature commonly referred as power law inflation (PLI)[65, 66, 67, 68]. Such potential was also studied in the context of M-theory [69] and RSII scenarios [70]. Due to the strong constraints imposed by Planck data [71] on PLI, we follow the generalization proposed in Refs.([72, 73]) named $\beta$-exponential inflation. In Ref.([72]), the authors introduced a class of potentials in a form
$V(\phi)=V_{0}\exp_{1-\beta}(-\lambda\phi/M_{pl})\;,$ (73)
$\;\;\;\;\;\;=V_{0}\left[1+\beta\left(-\lambda\phi/M_{pl}\right)\right]^{1/\beta}\;.$
In general, the function $\exp_{1-\beta}(f)=\left[1+\beta f\right]^{1/\beta}$ may be $01+\beta f>$ 0 or, otherwise, $\exp_{1-\beta}(f)=0$. Then, Eq.(73) should satisfy the identities $\exp_{1-\beta}(\ln_{1-\beta}(f))=f$ and $\ln_{1-\beta}(f)+\ln_{1-\beta}(g)=\ln_{1-\beta}(fg)-\beta\left[\ln_{1-\beta}(f)\ln_{1-\beta}(g)\right]$ for any $g<0$. The term $\ln_{1-\beta}(f)=(f^{\beta}-1)/\beta$ is referred as the generalized logarithm function. To our purposes, Eq.(72) can be generalized to Eq.(73) in a form
$V(\phi)=V_{0}\left[1-\beta\alpha_{0}\phi\right]^{1/\beta}\;.$ (74)
Assuming that the field starts rolling in a local minimum in such $\frac{\partial^{2}V(\phi)}{\partial^{2}\phi}=0$ for any $\beta\neq 1$, starting at $\phi=\phi_{\min}=1$, one obtains the condition $\beta=\frac{1}{\alpha_{0}}>0$. Thus, we can write Eq.(74) as
$V(\phi)=V_{0}\left[1-\phi\right]^{\alpha_{0}}\;.$ (75)
Then, one defines the pair of slow-roll parameters
$\varepsilon=\frac{1}{2\kappa}\left(\frac{V_{,\phi}}{V}\right)^{2}\;,$ (76)
$\eta=\frac{1}{\kappa}\frac{V_{,\phi\phi}}{V}\;.$ (77)
In order to submit the model to the scrutiny of observational data, they are expressed as
$\eta_{s}=1-6\varepsilon+2\eta\;,$ (78)
$r=16\varepsilon\;,$ (79)
where $\eta_{s}$ is the spectral tilt and $r$ is the tensor-to-scalar ratio. The relation between the parameters is given by
$\eta_{s}=1-\frac{\alpha_{0}(\alpha_{0}+2)}{(1-\phi_{\star})^{2}}\;,$ (80)
$r=\frac{8\alpha_{0}^{2}}{(1-\phi_{\star})^{2}}\;,$ (81)
where $\phi_{\star}$ is the field before the end of inflation given by
$\phi_{\star}=1-\alpha_{0}\sqrt{0.5+\frac{2}{\alpha_{0}}N}\;.$ (82)
The quantity $N$ denotes the number of e-folds before the end of inflation defined as
$N=\int^{\phi_{\star}}_{\phi_{end}}\frac{d\phi}{2\sqrt{\varepsilon}}=\frac{\phi_{\star}^{2}}{2\alpha_{0}}-\frac{\phi_{\star}}{\alpha_{0}}+\frac{1}{2\alpha_{0}}-\frac{\alpha_{0}}{4}\;.$ (83)
In the end of inflation, the field $\phi_{end}$ is calculated from the condition $\varepsilon(\phi_{end})\sim 1$, and one finds
$\phi_{end}=1-\frac{\alpha_{0}\sqrt{2}}{2}\;.$ (84)
Dataset ID Description Reference
sixdf_2011_bao 6dF Galaxy Survey [57]
sdss_dr7_mgs Seventh Data Release of SDSS Main Galaxy Sample (SDSS DR7 MGS) [58]
sdss_dr16_baoplus_lrg BOSS DR16 - Luminous Red Galaxies (LRG) [59]
sdss_dr16_baoplus_elg BOSS DR16 - Emission Line Galaxies (ELG) [74]
sdss_dr16_baoplus_qso BOSS DR16 - Quasars (QSO) [74]
sdss_dr16_baoplus_lyauto BOSS DR16 - Lyman-alpha forest auto-correlation [74]
sdss_dr16_baoplus_lyxqso BOSS DR16 - Lyman-alpha forest x Quasar cross-correlation [74]
Table 1: Datasets from SDSS employed in our analysis.
Once the slow-roll parameters as a function of $\phi$ are given by
$\varepsilon(\phi)=\frac{\alpha_{0}^{2}}{2(1-\phi)^{2}}\;,$ (85)
$\eta(\phi)=\frac{\alpha_{0}(\alpha_{0}-2)}{2(1-\phi_{\star})^{2}}\;.$ (86)
[…]
V Data and methodology
To delineate constraints on the free parameter of the theory framework under consideration, we will examine several datasets individually as well as in various combinations. Specifically, we will consider:
  • Utilizing the latest NPIPE Planck DR4 likelihoods [54, 55] of Planck 2018 legacy data release, our analysis incorporates Cosmic Microwave Background (CMB) measurements. These include the $-\ell$ Plik TT likelihood spanning the multipole range $30\leq\ell\leq 2508$, as well as TE and EE measurements within the multipole range $30\leq\ell\leq 1996$. Additionally, we incorporate $-\ell$ TT-only ($2\leq\ell\leq 29$) and EE-only ($2\leq\ell\leq 29$) likelihoods. Furthermore, our dataset encompasses CMB NPIPE Planck lensing power spectrum measurements, collectively referred to as the Planck dataset.
  • The B-mode polarization data from the BICEP2/Keck collaboration [56]. We refer to this data set as BK18.
  • We refer to this data set as LSS. We consider 6dF Galaxy Survey [57], the Seventh Data Release of SDSS Main Galaxy Sample (SDSS DR7 MGS) [58] and clustering measurements of the Extended Baryon Oscillation Spectroscopic Survey (eBOSS) associated with the SDSS’s Sixteenth Data Release [59]. This collection encompasses data from Luminous Red Galaxies (LRG), Emission Line Galaxies (ELG), Quasars (QSO), the Lyman-alpha forest auto-correlation (lyauto), and the Lyman-alpha forest x Quasar cross-correlation (lyxqso). Table 1 details the diverse BAO components considered in this work.
    • The Hubble distance at redshift $z$:
      $D_{H}(z)=\frac{c}{H(z)},$ (87)
      where $H(z)$ is the Hubble parameter.
    • The comoving angular diameter distance, $D_{M}(z)$, which also only depends on the expansion history:
[…]
    • The spherically-averaged BAO distance:
      $D_{V}(z)=r_{d}[zD_{M}^{2}(z)D_{H}(z)]^{1/3},$ (89)
      where $r_{d}$ is the BAO scale, of which in our analyzes we treat it as a derived parameter.
    Parameter Prior
    $\Omega_{b}h^{2}$ $\mathcal{U}[0.017,0.027]$
    $\Omega_{c}h^{2}$ $\mathcal{U}[0.09,0.15]$
    $\theta_{\mathrm{MC}}$ $\mathcal{U}[0.0103,0.0105]$
    $\tau_{\rm reio}$ $\mathcal{N}[0.065,0.0015]$
    $\log(10^{10}A_{\mathrm{s}})$ $\mathcal{U}[2.6,3.5]$
    $n_{s}$ $\mathcal{U}[0.9,1.1]$
    $10^{8}$ $\gamma_{s}$ $\mathcal{U}[-1,1]$
    Table 2: The cosmological parameters along with their respective priors employed in the parameter estimation analysis.
    For the growth measurements, the growth function $f$ can be expressed as a differential in the amplitude of linear matter fluctuations on a comoving scale of 8 $h^{-1}$ Mpc, $\sigma_{8}(z)$, in the form
    $f(z)=\frac{\partial\ln\sigma_{8}}{\partial\ln a}.$ (90)
    The RSD measurements provide constraints on the quantity $f(z)\sigma_{8}(z)$. The $\sigma_{8}(z)$ depends on the matter power spectrum, $P(k,z)$, which is calculated by default in the Boltzmann code. Both $f(z)$ and $\sigma_{8}(z)$ are sensitive to variations in the effective gravitational coupling and the light deflection parameter, which play a crucial role in our Poisson and lensing equations in MG models.
[…]
Parameter Planck+BK18 Planck+BK18+LSS
[…]
$n_{\mathrm{s}}$ $0.9646\pm 0.0031$ $0.9646^{+0.0057}_{-0.0060}$ $0.9648\pm 0.0030$ $0.9648^{+0.0056}_{-0.0060}$
$r$ $0.0142^{+0.0049}_{-0.012}$ $<0.0307$ $0.0139^{+0.0044}_{-0.013}$ $<0.0303$
$H_{0}$ $66.981\pm 0.078$ $66.98^{+0.15}_{-0.15}$ $67.001\pm 0.078$ $67.00^{+0.15}_{-0.15}$
$\Omega_{\mathrm{m}}$ $0.31794\pm 0.00074$ $0.3179^{+0.0014}_{-0.0015}$ $0.31776\pm 0.00074$ $0.3178^{+0.0014}_{-0.0014}$
$\mu-1$ $-0.01^{+0.17}_{-0.21}$ $-0.01^{+0.40}_{-0.34}$ $0.02\pm 0.15$ $0.02^{+0.30}_{-0.28}$
$\Sigma-1$ $-0.053^{+0.051}_{-0.042}$ $-0.053^{+0.093}_{-0.098}$ $-0.049^{+0.048}_{-0.042}$ $-0.049^{+0.083}_{-0.091}$
$10^{8}\gamma_{s}$ $-0.3^{+3.1}_{-3.8}$ $-0.3^{+7.2}_{-6.2}$ $0.3\pm 2.7$ $0.3^{+5.4}_{-5.0}$
Figure 2: Triangle plot contours with MG parameters $\mu-\Sigma$ and the $\gamma_{s}$ parameter of the $\beta$-model for the combined analyses of Planck+BK18 (red line) and Planck+BK18+LSS (blue line).
The joint likelihood analysis is conducted using MGCAMB-II [60] through the Cobaya [61] sampler, employing the $\mu-\Sigma$ parametrization defined as follows:
$\mu(a,k)=1+\mu_{0}\frac{\Omega_{DE}}{\Omega_{DE(0)}}\;,$ (91)
$\Sigma(a,k)=1+\Sigma_{0}\frac{\Omega_{DE}}{\Omega_{DE(0)}}\;.$ (92)
[…]
The priors on the baseline parameters utilized in our analysis are detailed in Table 2. In all runs, we ensure a Gelman-Rubin convergence criterion of $R-1<0.03$. In the subsequent section, we will unveil the outcomes of our Bayesian analysis and delve into their implications.
VI Results
We commence our analysis by scrutinizing the constraints derived exclusively from the joint analysis of Planck and BK18 datasets. The primary statistical findings concerning the cosmological parameters of interest are summarized in Table 3.
We commence our analysis of the results by examining the parameters that indicate deviations from the standard $\Lambda$CDM cosmology. Figure (2) illustrates the parameter space for $\gamma_{s}$, $\mu$, and $\Sigma$.
Figure 3: The contours in the $n_{s}-r$ plane, delineating 68% and 95% confidence levels (C.L.), represent the combined analyses of Planck+BK18 (grey dashed line contour) and Planck+BK18+LSS (blue line contour) using the $\beta$-model. The red contour illustrates the joint fitting of Planck+BK18+LSS extracted from publicly available CosmoMC chains [56].
[…]
Considering the scalar spectral index $n_{s}$, we obtained a close value of $n_{s}=0.9641\pm 0.0031$ with Planck+BK18+LSS data at $68\%$ CL to based-$\Lambda$CDM Planck TT,TE,EE+lowE+lensing with $n_{s}=0.9649\pm 0.0042$ at $68\%$ CL and $n_{s}=0.9665\pm 0.0038$ at $68\%$ CL when considering Planck TT,TE,EE+lowE+lensing+LSS. On the other hand, the small differences between the $n_{s}$ value seem to affect the value of $H_{0}$ which is roughly $0.5\%$ lower than the Planck-$\Lambda$CDM case. From Planck+BK18+LSS dateset, in the analysis of slow-roll parameters, the model prefers the case where $\alpha_{0}=0.13$ gives $w=-0.9944$ as shown in Figure (1) for the number of e-folds $N=50$ and $N=60$. Most importantly, the model does not require a larger number of e-folds to provide a tighter constraint on $r$. We also verified that the inclusion of LSS data does not significantly change the value of $n_{s}$ but influences the $H_{0}$ values, which may be improved with upcoming new constraints on the reionization optical depth whose uncertainties may provoke a large impact on fundamental cosmological parameters such as $n_{s}$ [76].
VII Final Remarks
In this paper, we have derived the gravitational equations within four dimensions by inducing them from a five-dimensional bulk, employing Nash’s embeddings framework and incorporating them into a well-established $\mu$ $\Sigma$ MG representation. From the analysis of the slow-roll conditions, we have obtained a PLI model that was generalized to the $\beta$-exponential inflation [72, 73] that suggest $w=-0.9944$ for the number of e-folds in the range $N=50$ and $N=60$. Interestingly, this model also sits within the previous reasonable expected number e-folds compatible with the tighter restriction on $r$, which is important to maintain the window of solving the horizon problem. This enables us to assess and investigate the impact of the model on linear perturbations in the CMB data. Apart from the values of the tensor-to-ratio parameter, our primary analyses revealed similar predictions to the ones of the $\Lambda$CDM model. Additionally, we have quantified the model’s predictions concerning inflationary dynamics. By utilizing data from CMB-PR4, BICEP/Keck Array 2018, and certain LSS measurements, we have established an tighter upper limit of $r<0.0303$ at a 95% C.L. As compared with $\Lambda$CDM model, such apparent improvement of tensor-to-ratio parameter and a similar but lower value of the scalar spectral index $n_{s}$ values may suggest a glimpse of a modified gravity signature which may be improved in future experiments.
Acknowledgements.
AJSC thanks the financial support from the Conselho Nacional de Desenvolvimento Científico e Tecnológico (CNPq, National Council for Scientific and Technological Development) for partial financial support under the project No. 305881/2022-1, and The Fundação da Universidade Federal do Paraná (FUNPAR, Paraná Federal University Foundation) by public notice 04/2023-Pesquisa/PRPPG/UFPR- for partial financial support under process nº 23075.019406/2023-92. R.C.N thanks the financial support from CNPq under the project No. 304306/2022-3, and the Fundação de Amparo à pesquisa do Estado do RS (FAPERGS, Research Support Foundation of the State of RS) for partial financial support under the project No. 23/2551-0000848-3.